High-power experimental engines are being developed by the Stevens Motor Company for use in its new sports coupe. The engineers have calculated the maximum horsepower for the engine to be 830HP. Twenty five engines are randomly selected for horsepower testing. The sample has an average maximum HP of 880 with a standard deviation of 55HP. Assume the population is normally distributed.

Calculate a confidence interval for the average maximum HP for the experimental engine. Use a significance level of α=0.01. Round your answers to two decimal places.
Answer #1

Confidence Interval

\bar{X} \pm t_{\alpha /2, n-1} S/\sqrt{n}

t_{\alpha /2, n-1} = t_{ 0.01 /2, 25- 1 } = 2.797

880 \pm t_{\ 0.01/2, 25 -1} * 55/\sqrt{ 25}

Lower Limit = 880 - t_{\ 0.01/2, 25 -1}55/\sqrt{ 25}

Lower Limit = 849.23

Upper Limit = 880 + t_{\ 0.01/2, 25 -1}55/\sqrt{ 25}

Upper Limit = 910.77

99% Confidence interval is ( 849.23 , 910.77 )